OpenStudy (anonymous):

Name the property: 2 + (3 + 4) = (2 + 3) + 4 Distributive Propertty Associative Property of Addition Commutative Property of Addition Additive Identity

OpenStudy (anonymous):

This is really easy once you get it, but you have to know the properties.

OpenStudy (anonymous):

The DISTRIBUTIVE PROPERTY: 5 (x + 2) 5x + 10 The ASSOCIATIVE PROPERTY OF ADDITION: (2 + 3) + 4 = 2 + (3 + 4) So the answer to your problem is "The Associative Property of Addition" because it is the same format.

OpenStudy (anonymous):

Associative Property of Addition

OpenStudy (anonymous):

ADDITIVE IDENTITY: 5 + 0 = 5 (it's called that because the number stays the same) COMMUTATIVE PROPERTY OF ADDITION: 2 + 3 = 3 + 2